What is the product of  x(x-a)(x-b)...................................(x-z)

Grade:12

2 Answers

Arun
4 years ago
Dear Student
 
Answer will be 0
 
As in somewhere (x-x) will come.
And (x-x) = 0
Hence 
0 * something = 0
 
Hope it helps
Vikas TU
4 years ago
Dear student 
We have to find,
(X-A)(X-B)(X-C)....(X-Y)(X-Z)
= (X-A)(X-B)(X-C)(X-D)....(X-W)(X-X)(X-Y)(X-Z)
= (X-A)(X-B)(X-C)(X-D)....(X-W)(0)(X-Y)(X-Z) (since X-X = 0)
= 0 [Product of any term multiplied with zero always results in zero].
Hence the answer is zero.